# Kalman Filter Enhanced Group Relative Policy Optimization for Language Model Reasoning

### v4

@@ Appendix B Model Performance on More Base-models
To evaluate the generalization with larger models, we also conducted experiments with Llama3.2-3B-Instruct model as shown in the Fig. 5 and Fig. 6. With the Llama3.2-3B-Instruct and Qwen2.5-7B-Instruct models, the training returns also demonstrate that our proposed KRPO method consistently achieves higher rewards than the GRPO baseline, confirming the effectiveness of our approach.
Figure 5: The training return curves of Llama3.2-3B-Instruct model on normal level of Arithmetic data.
Figure 6: The training return curves of Qwen2.5-7B-Instruct model on normal level of Arithmetic data.
@@ Appendix D More Comparisons with Other Models
With respect to the baseline comparison, we formulated the Arithmetic-hard data into preferred and unpreferred pairs to enable optimization with Direct Preference Optimization (DPO) (Rafailov et al., 2023). The experimental results shown in the Tab. 5 indicate that KRPO remains the best-performing method for this task. This is due to its improved reward baseline and uncertainty estimation, which leads to more reliable advantage signals during optimization.
To have a more comprehensive comparison, we also conduct experiments to test the actual performance of fixed baselines. We found that the “fixed value” baseline raises a practical problem of what value to use. We compute the advantages within each group following the GRPO setup. As noted in the implementation details, we use a group size of 12, which means the rewards range from 0 to 12. Therefore, we chose 0 (without considering baseline), 6 (the midpoint) and 3 (the first quartile) as fixed-value baselines. If set to 0, the algorithm degenerates into a vanilla policy gradient without considering advantages. We tested this case and found that while learning is still possible (the accuracy is only 11.901% as shown in the table), the optimization is less stable. If we set the baseline to 3 or 6, the optimization collapses entirely (approx. 0.000% accuracy). In the table, “GRPO-fixed-b0” denotes using 0 as the fixed baseline value, while “GRPO-fixed-b3” and “GRPO-fixed-b6” represent using 3 and 6 as fixed baseline values, respectively. This aligns with the understanding in the policy gradient literature of Sutton et al. (Sutton et al., 1999), where an effective baseline should approximate the expected return conditioned on the current policy. A fixed baseline cannot adapt to varying expected returns across states or tasks, resulting in it cannot fulfill the purpose of reducing the variance of policy gradient. So, such an approach is generally unsuitable for practical applications in most reinforcement learning settings.
@@ Appendix E Model Performance with Different Group Size
Therefore, we also conducted experiments on model performance with different group sizes (= {5, 8, 10, 15}) as shown in Fig. 7(a). In the figure, “gx” denotes with group size x (e.g. g5 means with group size 5). Across various group size settings, the proposed KRPO consistently outperformed the baseline GRPO, demonstrating its robustness and effectiveness. Interestingly, as the group size varies from 5 to 15, the performance of both KRPO and GRPO improves; however, a large performance gap always remains between them, with KRPO maintaining a clear advantage. Interestingly, we empirically observe that as the number of group samples are very few, the performance gap between KRPO and its GRPO baseline widens, indicating that KRPO is particularly effective at handling extremely noisy rewards when only very sparse group samples are available.
@@ Appendix F Model Performance with Different Random Seeds
(a) Model performance with different group sizes.
(b) Model performance with different random seeds.
@@ Appendix G Time Efficiency
Appendix H More Examples of Case Study
@@
Arithmetic {“question”: “What is -( -985 - -604 )?”, “answer”: “381”} The calculation process involves: -( (-985 - (-604) ) ) = -( (-985 + 604 ) ) = -( -381 ) = 381
Arithmetic {“question”: “What is 9147 + -1895 + -449?”, “answer”: “6803”} Calculating the expression: 9147 - 1895 = 7252, then 7252 - 449 = 6803
Arithmetic {“question”: “( -74 * -25 ) + -34 = ?”, “answer”: “1816”} First, -74 * -25 = 1850, then 1850 + -34 = 1816
Arithmetic {“question”: “Solve -( 427 - 202 ) + 514”, “answer”: “289”} Let’s calculate this step by step: -( 427 - 202 ) + 514 = -( 225 ) + 514, -( 225 ) = -225, -225 + 514 = 289
OpenMath {“type”: “Prealgebra”, “question”: “Add $81.76$ to $34.587$ and round your answer to the nearest tenth.”} To add 81.76 and 34.587, we first add the numbers: 81.76 + 34.587 = 116.347. Then, we round to the nearest tenth, which gives us 116.3.
OpenMath {“type”: “Prealgebra”, “question”: “What is the greatest common factor of 32 and 48”, “expected_answer”: “16”} To find the greatest common factor (GCF) of 32 and 48, we need to list the factors of each number. The factors of 32 are: 1, 2, 4, 8, 16, 32. The factors of 48 are: 1, 2, 3, 4, 6, 8, 12, 16, 24, 48. The greatest common factor is the largest number that appears in both lists, which is 16.
OpenMath {“type”: “Counting & Probability”, “question”: “Compute $\dbinom{7}{4}$.”, “expected_answer”: “35”} To compute $\dbinom{7}{4}$, we use the formula for combinations: $\dbinom{n}{r}=\dfrac{n!}{r!(n-r)!}$. In this case, $n=7$ and $r=4$. Plugging these values into the formula, we get: $\dbinom{7}{4}=\dfrac{7!}{4!(7-4)!}=\dfrac{7!}{4!3!}=\dfrac{7\times 6\times 5}{3\times 2\times 1}=35$.
@@ Appendix I The Analysis of Leading Statement for Language Model Reasoning
From Tab. 1 and Tab. 6 we find that compared with OpenMath-Instruct data, the Arithmetic data questions are not harder, but the model receives much lower results, especially in the hard arithmetic questions. So we hypothesize that the leading statement might be a reason, as there are leading statements in OpenMath-Instruct data, but not in Arithmetic data. So we add leading statements, “Please calculate the following expression: ” to each arithmetic question.
@@
Figure 8: The curves of (a) KL divergence; (b) normalized gradients of GRPO and the proposed KRPO; and (c) the Kalman filter gradually stabilizes the reward baseline to a certain level.
@@
We provide an example to show the stabilization of the baseline from the Kalman Filter in Fig. 8(c). With 12 rewards between [0,1], the Kalman filter gradually stabilizes the reward baseline.

### v5

@@ Appendix B Model Performance on More Base-models
To evaluate the generalization with larger models, we also conducted experiments with Llama3.2-3B-Instruct model as shown in the Fig. 6(a) and Fig. 6(b). With the Llama3.2-3B-Instruct and Qwen2.5-7B-Instruct models, the training returns also demonstrate that our proposed KRPO method consistently achieves higher rewards than the GRPO baseline, confirming the effectiveness of our approach.
(a) Llama3.2-3B-Instruct
(b) Qwen2.5-7B-Instruct
Figure 6: The training return curves of (a) Llama3.2-3B-Instruct model and (b) Qwen2.5-7B-Instruct model on normal level of Arithmetic data.
@@ Appendix D More Comparisons with Other Models
With respect to the baseline comparison, we formulated the Arithmetic-hard data into preferred and unpreferred pairs to enable optimization with Direct Preference Optimization (DPO) [18]. The experimental results shown in the Tab. 4 indicate that KRPO remains the best-performing method for this task. This is due to its improved reward baseline and uncertainty estimation, which leads to more reliable advantage signals during optimization.
To have a more comprehensive comparison, we also conduct experiments to test the actual performance of fixed baselines. We found that the “fixed value" baseline raises a practical problem of what value to use. We compute the advantages within each group following the GRPO setup. As noted in the implementation details, we use a group size of 12, which means the rewards range from 0 to 12. Therefore, we chose 0 (without considering baseline), 6 (the midpoint) and 3 (the first quartile) as fixed-value baselines. If set to 0, the algorithm degenerates into a vanilla policy gradient without considering advantages. We tested this case and found that while learning is still possible (the accuracy is only 11.901% as shown in the table), the optimization is less stable. If we set the baseline to 3 or 6, the optimization collapses entirely (approx. 0.000% accuracy). In the table, “GRPO-fixed-b0” denotes using 0 as the fixed baseline value, while “GRPO-fixed-b3” and “GRPO-fixed-b6” represent using 3 and 6 as fixed baseline values, respectively. This aligns with the understanding in the policy gradient literature of Sutton et al. [22], where an effective baseline should approximate the expected return conditioned on the current policy. A fixed baseline cannot adapt to varying expected returns across states or tasks, resulting in it cannot fulfill the purpose of reducing the variance of policy gradient. So, such an approach is generally unsuitable for practical applications in most reinforcement learning settings.
@@ Appendix E Model Performance with Different Group Size
Therefore, we also conducted experiments on model performance with different group sizes (= {5, 8, 10, 15}) as shown in Fig. 4(a). In the figure, “gx” denotes with group size x (e.g. g5 means with group size 5). Across various group size settings, the proposed KRPO consistently outperformed the baseline GRPO, demonstrating its robustness and effectiveness. As the group size varies from 5 to 15, the performance of both KRPO and GRPO improves; however, a large performance gap always remains between them, with KRPO maintaining a clear advantage. Interestingly, we empirically observe that as the number of group samples are very few, the performance gap between KRPO and its GRPO baseline widens, indicating that KRPO is particularly effective at handling noisy rewards when only very sparse group samples are available.
@@ Appendix G Time Efficiency
Appendix H Case Study
Question {“type": “Algebra", “question": “If $7^{4x}=343$, what is the value of $7^{4x-3}$", “expected_answer": “1"}
Model Thinking Process
We also provide a case study from OpenMath data as Tab. 5 to show the improved reasoning ability of the KRPO model over the GRPO baseline. From the case, we observe that GRPO can figure out the prior steps, e.g. “This implies $4x=3$, so $x=\frac{3}{4}$”, but it stops here and derives the wrong answer 343. In contrast, the proposed KRPO can get the correct answer 1. For this question, the KRPO can get 12 corrects out of 12 trials, while the GRPO can have 3 corrects out of 12. This qualitative example is consistent with the hypothesis that KRPO provides a more useful training signal than GRPO on this task, although a single case study does not isolate the causal mechanism.
@@ Appendix H Case Study
Arithmetic {“question": “What is -( -985 - -604 )?", “answer": “381"} The calculation process involves: -( (-985 - (-604) ) ) = -( (-985 + 604 ) ) = -( -381 ) = 381
Arithmetic {“question": “What is 9147 + -1895 + -449?", “answer": “6803"} Calculating the expression: 9147 - 1895 = 7252, then 7252 - 449 = 6803
Arithmetic {“question": “( -74 * -25 ) + -34 = ?", “answer": “1816"} First, -74 * -25 = 1850, then 1850 + -34 = 1816
Arithmetic {“question": “Solve -( 427 - 202 ) + 514", “answer": “289"} Let’s calculate this step by step: -( 427 - 202 ) + 514 = -( 225 ) + 514, -( 225 ) = -225, -225 + 514 = 289
OpenMath {“type": “Prealgebra", “question": “Add $81.76$ to $34.587$ and round your answer to the nearest tenth."} To add 81.76 and 34.587, we first add the numbers: 81.76 + 34.587 = 116.347. Then, we round to the nearest tenth, which gives us 116.3.
OpenMath {“type": “Prealgebra", “question": “What is the greatest common factor of 32 and 48", “expected_answer": “16"} To find the greatest common factor (GCF) of 32 and 48, we need to list the factors of each number. The factors of 32 are: 1, 2, 4, 8, 16, 32. The factors of 48 are: 1, 2, 3, 4, 6, 8, 12, 16, 24, 48. The greatest common factor is the largest number that appears in both lists, which is 16.
OpenMath {“type": “Counting & Probability", “question": “Compute $\dbinom{7}{4}$.", “expected_answer": “35"} To compute $\dbinom{7}{4}$, we use the formula for combinations: $\dbinom{n}{r}=\dfrac{n!}{r!(n-r)!}$. In this case, $n=7$ and $r=4$. Plugging these values into the formula, we get: $\dbinom{7}{4}=\dfrac{7!}{4!(7-4)!}=\dfrac{7!}{4!3!}=\dfrac{7\times 6\times 5}{3\times 2\times 1}=35$.
@@ Appendix I The Analysis of Leading Statement for Language Model Reasoning
From Tab. 1 and Tab. 6 we find that compared with OpenMath-Instruct data, the Arithmetic data questions are not harder, but the model receives much lower results, especially in the hard arithmetic questions. So we hypothesize that the leading statement might be a reason, as there are leading statements in OpenMath-Instruct data, but not in Arithmetic data. So we add leading statements, “Please calculate the following expression: " to each arithmetic question.
@@
Figure 7: The curves of (a) KL divergence; (b) normalized gradients of GRPO and the proposed KRPO; and (c) an illustration shows the Kalman filter gradually stabilizes the reward baseline to a certain level.
@@
We provide an illustration to show the stabilization of the baseline from the Kalman Filter in Fig. 7(c). With 12 rewards between [0,1], the Kalman filter gradually stabilizes the reward baseline.